# Oblique boundary value problems for degenerate Hessian quotient type equations

### v1

@@ Abstract.
In this article, we study oblique boundary value problems for degenerate Hessian quotient type equations in bounded domains without any geometric conditions. By establishing the a priori estimates, we show the existence and uniqueness of $C^{1,1}$ admissible solutions.
@@
This research was supported by funds from the National Natural Science Foundation of China No. 11971157.
@@ 1. Introduction
In this article, we consider oblique boundary value problems for degenerate Hessian quotient type equations
(1.1) $&\frac{\sigma_{k}(\gamma\triangle uI-D^{2}u)}{\sigma_{l}(\gamma\triangle uI-D^{2}u)}=f(x,u),\,\,\,\text{in}\,\;\Omega,\\ &G(x,u,Du)=0,\,\,\,\,\quad\quad\quad\quad\quad\text{on}\,\;\partial\Omega,\end{aligned}\right.$
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$, $\gamma\geq 1$, $0\leq l<k\leq n$, $Du$ and $D^{2}u$ denote the gradient vector and the Hessian matrix of the solution $u$. For convenience, let $x,z,p$ be points in $\Omega,\mathbb{R},\mathbb{R}^{n}$, respectively. Set $f(x,z)$ is a non-negative smooth function defined on $\Omega\times\mathbb{R}$. The boundary value problem $G(x,z,p)\in C^{1}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ is said to be oblique with respect to $u$, if
(1.2) $G_{p}\cdot\nu\geq\beta_{0},\,\,\text{for all}\,\;(x,z,p)\in\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n},$
where $\nu$ is the unit inner normal vector field on $\partial\Omega$ and $\beta_{0}$ is a positive constant. If $G_{p}\cdot\nu>0$ on all of $\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n}$, we will simply refer to $G$ as oblique. As in Jiang-Trudinger-Xiang [24], we will assume that the function $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ is concave in $p$, that is $G_{pp}\leq 0$ on $\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n}$. This includes the quasilinear case, when $G_{pp}=0,$
(1.3) $G(x,z,p)=\beta(x,z)\cdot p-\varphi(x,z)=0,\quad\text{on}\;\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n},$
where $\beta=G_{p}$ and $\varphi$ are defined on $\partial\Omega\times\mathbb{R}$. In this article, we focus on the semilinear case,
(1.4) $u_{\beta}=\varphi(x,u),\quad\text{on}\ \partial\Omega\times\mathbb{R},$
where $\beta\cdot\nu>0$. Especially, we need $\varphi_{u}\geq\gamma_{0}>0$ which is only used to prove upper bounds of the admissible solution.
@@
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, we set the elementary symmetric function for $k=1,2,\cdots,n$
@@
[…]
@@
(1.5) $\frac{\sigma_{k}(\gamma\triangle uI-D^{2}u)}{\sigma_{l}(\gamma\triangle uI-D^{2}u)}=f$
@@
For non-degenerate cases, Dirichlet problems for basic Hessian equations have been well studied by Caffarelli-Nirenberg-Spruck [1], Ivochkina [19] and Trudinger [39]. Some works about Neumann problems for basic Hessian equations can be found in [4, 33, 36, 38]. Dong-Wei [7] have established the a priori estimates to equation (1.5) when $\lambda(\triangle uI-D^{2}u)(x)\in\Gamma_{k+1}$. Then Chen-Dong-Han [5] showed that the existence and uniqueness of the admissible solution to equation (1.5) with $0\leq l<k<n$ when $\lambda(\triangle uI-D^{2}u)(x)\in\Gamma_{k}$.
Let’s review some research about degenerate fully nonlinear equations. The a priori $C^{1,1}$ estimates play an important role in the proof of $C^{1,1}$ solutions in degenerate problems. Guan [12] established $C^{1,1}$ estimates to Dirichlet problems of degenerate Monge-Ampère equations for homogenous boundary condition with $f^{\frac{1}{n}}\in C^{1,1}(\overline{\Omega})$, and Guan-Trudinger-Wang [13] for non-homogenous boundary condition with $f^{\frac{1}{n-1}}\in C^{1,1}(\overline{\Omega})$. Dong [6] obtained $C^{1,1}$ estimates to Dirichlet problems of degenerate Hessian equations for homogenous boundary condition with $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$ and $Df^{\frac{1}{k-1}}\leq Cf^{\frac{1}{2(k-1)}}$. Recently, Jiao-Wang [20] solved $C^{1,1}$ estimates to convex solutions of Dirichlet problems for degenerate Hessian equations with non-homogenous boundary conditions defined in the strictly convex bounded domain with $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$. If $f^{\frac{1}{k}}\in C^{1,1}(\overline{\Omega})$ for non-homogenous boundary problems, $C^{1,1}$ estimates have been obtained by Krylov [26, 27, 28, 29]. Wang [47] gave an example to show that the condition $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$ is sharp for degenerate Hessian equations. The existence of $C^{1,1}$ solutions to Neumann problems for degenerate basic Hessian quotient equations with $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega})$ was proved by Mei [34]. For more reference, the reader may see [2, 11, 48] and the reference therein.
@@
Therefore, it is of natural interest to consider existence results for the oblique boundary value problem. The oblique boundary value problem for Hessian type equations arises naturally in the theory of fully nonlinear elliptic equations, through its applications in conformal geometry, optimal transportation, and geometric optics. Wang [45] derived the existence of solutions for two dimensional Monge-Ampère equations of the oblique boundary value problem and considered the generalized solutions for general dimension. The oblique boundary value problem of linear and quasilinear elliptic equations can be seen in the book [30]. Lieberman-Trudinger [31] considered the nonlinear oblique boundary value problem for quasilinear and fully nonlinear uniformly elliptic partial differential equations under some structure conditions. Urbas [42, 43, 44] studied Monge-Ampère equations, two dimensional Hessian equations and curvature equations of the oblique boundary value problem. For augmented Hessian equations with the oblique boundary value problem
@@
Inspired by the above works, we naturally consider the oblique boundary value problem (1.1) for degenerate Hessian quotient type equations and get the following main theorem. In order to obtain the existence, we need to assume the existence of an admissible subsolution $\underline{u}$ to problem (1.1) satisfying
(1.6) $&\frac{\sigma_{k}(\gamma\triangle\underline{u}I-D^{2}\underline{u})}{\sigma_{l}(\gamma\triangle\underline{u}I-D^{2}\underline{u})}\geq f(x,\underline{u}),\,\,\,\text{in}\,\;\Omega,\\ &G(x,\underline{u},D\underline{u})=0,\,\,\,\,\quad\quad\quad\quad\quad\text{on}\,\;\partial\Omega.\end{aligned}\right.$
Then we have the following existence result for $\gamma\geq 1$ and $0\leq l<k<n$.
@@
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{4}$ domain $\Omega\subset\mathbb{R}^{n}$, $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$ and $f_{z}\geq 0$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the oblique boundary condition (1.4). Assume that $\underline{u}$ is an admissible subsolution of problem (1.1). Then there exists a unique solution $u\in C^{1,1}(\overline{\Omega})$ of problem (1.1) satisfying
@@
where the positive constant $C$ depends on $n,\ k,\ l,\ \gamma_{0},\ \gamma,\ \overline{\Omega},$ $\underline{u}$, $G$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$.
@@
The assumption of the admissible subsolution and $f_{z}\geq 0$ are only used to prove maximum modulus estimates. Furtherly, When $\gamma>1$ in problem (1.1), the above existence result still holds for $k=n$.
@@
When the matrix $\left\{\gamma\triangle uI-D^{2}u\right\}$ ($\gamma\geq 1$) is replaced by $\left\{\gamma\triangle uI+D^{2}u\right\}$ ($\gamma>0$) in problem (1.1), the existence theorem still holds for the problem
(1.8) $&\frac{\sigma_{k}(\gamma\triangle uI+D^{2}u)}{\sigma_{l}(\gamma\triangle uI+D^{2}u)}=f(x,u),\,\,\,\text{in}\,\;\Omega,\\ &G(x,u,Du)=0,\quad\quad\quad\quad\quad\ \ \text{on}\,\;\partial\Omega,\end{aligned}\right.$
@@
The rest of the paper is organized as follows. In Section 2, some preliminaries are provided. Gradient estimates are established in Section 3. Section 4 are devoted to second order estimates and the proof of Theorem 1.3.
@@ 2. Preliminaries
In this section, we recall some basic properties of Hessian quotient type equations, which are well-known and can be found in [3, 5, 7].
Firstly, we define some notations for convenience. Let $U=\gamma\triangle uI-D^{2}u,U_{ij}=\gamma\triangle u\delta_{ij}-u_{ij}$, $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ and $\eta=(\eta_{1},\eta_{2},\cdots,\eta_{n})$ are eigenvalue vectors of matrix $D^{2}u$ and $U$, respectively. Also, we set
(2.1) $F(u_{ij})=\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})},\quad F^{ij}=\frac{\partial F}{\partial u_{ij}},\quad F^{ij,rs}=\frac{\partial^{2}F}{\partial u_{ij}\partial u_{rs}},$
and
(2.2) $\widetilde{F}(u_{ij})=\left[\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})}\right]^{\frac{1}{k-l}},\quad\widetilde{F}^{ij}=\frac{\partial\widetilde{F}}{\partial u_{ij}},\quad\widetilde{F}^{ij,rs}=\frac{\partial^{2}\widetilde{F}}{\partial u_{ij}\partial u_{rs}},$
@@
Differentiating equation (2.3) with respect to vector fields $\xi=(\xi^{1},\xi^{2},\cdots,\xi^{n})$ twice with $\xi^{i}\in C^{2}(\overline{\Omega})$, $i=1,2,\cdots,n.$, we have
(2.4) $F^{ij}u_{ij\xi}=\partial_{\xi}f:=f_{x_{\xi}}+f_{z}u_{\xi},$
(2.5) $\widetilde{F}^{ij}u_{ij\xi}=\partial_{\xi}\widetilde{f}:=\widetilde{f}_{x_{\xi}}+\widetilde{f}_{z}u_{\xi},$
@@
(2.6) $F^{ij,rs}u_{ij\xi}u_{rs\xi}+F^{ij}u_{ij\xi\xi}=\partial_{\xi\xi}f:=f_{x_{\xi}x_{\xi}}+2f_{x_{\xi}z}u_{\xi}+f_{zz}u_{\xi}^{2}+f_{z}u_{\xi\xi},$
(2.7) $\widetilde{F}^{ij,rs}u_{ij\xi}u_{rs\xi}+\widetilde{F}^{ij}u_{ij\xi\xi}=\partial_{\xi\xi}\widetilde{f}:=\widetilde{f}_{x_{\xi}x_{\xi}}+2\widetilde{f}_{x_{\xi}z}u_{\xi}+\widetilde{f}_{zz}u_{\xi}^{2}+\widetilde{f}_{z}u_{\xi\xi}.$
Next, we will introduce some properties of elementary symmetric functions.
Definition 2.1.
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$.
(1) $\sigma_{l}(\lambda|i)=\sigma_{l}(\lambda)|_{\lambda_{i}=0},$ $\sigma_{k-1}(\lambda|i)$ is also denoted by $\sigma_{k}^{ii}=\frac{\partial\sigma_{k}}{\partial\lambda_{i}},$
(2) $\sigma_{l}(\lambda|ij)=\sigma_{l}(\lambda)|_{\lambda_{i}=\lambda_{j}=0},$ $\sigma_{k-2}(\lambda|ij)$ is also denoted by $\sigma_{k}^{ii,jj}=\frac{\partial^{2}\sigma_{k}}{\partial\lambda_{i}\partial\lambda_{j}}.$
@@
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\Gamma_{k}$ and $1\leq k\leq n$, suppose that $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n},$ then
@@
where $C(n,k)$ is a positive constant only depending on $n$ and $k$.
The proof of Lemma 2.2 and the following generalized Newton-MacLaurin inequality can be respectively found in [32] and [37].
@@
If $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\widetilde{\Gamma}_{k}$, then $\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]^{\frac{1}{k-l}}(0\leq l<k\leq n)$ are concave with respect to $\lambda$. Hence, for any $\xi\in\mathbb{R}^{n}$, we have
$\sum_{i,j,r,s}\frac{\partial^{2}\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial u_{ij}\partial u_{rs}}\leq\left(1-\frac{1}{k-l}\right)\frac{\left[\sum\limits_{i,j}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial u_{ij}}\xi_{ij}\right]^{2}}{\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}},$
where $\widetilde{\Gamma}_{k}$ is an open convex symmetric cone in $\mathbb{R}^{n}$, given by
@@
Suppose $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\widetilde{\Gamma}_{k}$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, $\eta_{i}=\gamma\sum_{j=1}^{n}\lambda_{j}-\lambda_{i}$, for $0\leq l<k<n$ we obtain
@@
(3) $\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{i}}\geq\frac{2\gamma-1}{n\gamma-1}C(n,k,l)\sum_{j=1}^{n}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{j}}\geq\frac{2\gamma-1}{n\gamma-1}C(n,k,l)f^{1-\frac{1}{k-l}}\;(1\leq i\leq n).$
@@
When the matrix $\left\{\gamma\triangle uI-D^{2}u\right\}$ ($\gamma\geq 1$) is replaced by $\left\{\gamma\triangle uI+D^{2}u\right\}$ ($\gamma>0$), the similar Lemma 2.4 and Lemma 2.5 still hold for $0\leq l<k\leq n$. Furthermore, gradient estimates and second order estimates can be obtained and we omit details of the proof.
3. Gradient estimates
In this section, we prove gradient estimates for admissible solutions of problem (1.1).

### v2

# The oblique boundary value problem for degenerate Hessian quotient type equations

@@ Abstract.
In this paper, we investigate the oblique boundary value problem for degenerate Hessian quotient type equations in a smooth bounded domain. Without imposing any geometric restrictions on the domain, we establish the a priori estimates and derive the existence and uniqueness of admissible $C^{1,1}$ solutions under the condition $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$.
@@
This research was supported by funds from the National Natural Science Foundation of China No. 11971157, No. 12426532.
@@ 1. Introduction
This article is devoted to the study of oblique boundary value problems for degenerate Hessian quotient type equations
(1.1) $&\frac{\sigma_{k}}{\sigma_{l}}(\gamma\triangle uI-D^{2}u)=f(x,u),\,\,\,\text{in}\,\;\Omega,\\ &G(x,u,Du)=0,\,\,\,\,\quad\quad\quad\quad\quad\text{on}\,\;\partial\Omega,\end{aligned}\right.$
where $\Omega$ is a smooth bounded domain in $\mathbb{R}^{n}$, $\gamma\geq 1$, $0\leq l<k\leq n$, $Du$ and $D^{2}u$ denote the gradient vector and the Hessian matrix of the solution $u$. Here $f(x,u)$ is a non-negative function defined on $\overline{\Omega}\times\mathbb{R}$.
For convenience, let $x,z,p$ be points in $\Omega,\mathbb{R},\mathbb{R}^{n}$, respectively. The boundary value condition $G\in C^{1}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ is said to be strictly oblique with respect to $u$, if
(1.2) $G_{p}(x,u,Du)\cdot\nu\geq\beta_{0}>0,\,\,\text{for all}\,\;(x,u,Du)\in\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n},$
where $\nu$ is the unit inner normal vector field on $\partial\Omega$ and $\beta_{0}$ is a positive constant. If $G_{p}\cdot\nu>0$ on all of $\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n}$, we will simply refer to $G$ as oblique. As in Jiang-Trudinger-Xiang [27], we assume $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ is concave in $p$ if $G_{pp}\leq 0$ on $\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n}$. This includes the quasilinear oblique case, when $G_{pp}=0$
(1.3) $G=\beta(x,u)\cdot p-\varphi(x,u)=0,\quad\text{on}\;\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n},$
where $\beta(x,u)=G_{p}$, $p=Du$ and $\varphi(x,u)$ are defined on $\partial\Omega\times\mathbb{R}$. In this paper, we focus on the semilinear oblique boundary condition where $\beta$ in (1.3) is independent of $u$ with $p=Du$ s.t.
(1.4) $G=\beta(x)\cdot p-\varphi(x,u)=0,\quad\text{on}\ \partial\Omega\times\mathbb{R},$
which is equivalent to $u_{\beta}=\varphi(x,u)$. The standard example of (1.4) is the semilinear Neumann boundary condition, where $\beta=\nu$ on $\partial\Omega$.
@@
Assume $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, the elementary symmetric function for $k=1,2,\cdots,n$ is defined by
@@
We say $u\in C^{2}(\Omega)$ is $k$-admissible if $\lambda(\gamma\triangle uI-D^{2}u)(x)\in\Gamma_{k}$ for any $x\in\Omega$, $\mathring{\text{a}}$rding’s cone $\Gamma_{k}$ is
@@
(1.5) $\frac{\sigma_{k}}{\sigma_{l}}(\gamma\triangle uI-D^{2}u)=f$
@@
For non-degenerate cases, the Dirichlet problem for basic Hessian equations has been well studied by Caffarelli-Nirenberg-Spruck [1], Ivochkina [22] and Trudinger [38]. Some works about the Neumann problem for basic Hessian equations can be found in [4, 32, 35, 37]. Dong-Wei [7] have established the a priori estimates to equation (1.5) when $\gamma=1$, $\lambda(\triangle uI-D^{2}u)(x)\in\Gamma_{k+1}$. Then Chen-Dong-Han [5] showed that the existence and uniqueness of the $k$-admissible solution to equation (1.5) with $\gamma=1$, $0\leq l<k<n$.
We turn our attention to the degenerate case, where such estimates become more delicate and require refined techniques. The a priori $C^{1,1}$ estimates remain pivotal in establishing the regularity of solutions for these degenerate problems. For degenerate Monge-Ampère equations, foundational work by Guan-Li [15, 16] on the degenerate Weyl problem and Gauss curvature measure established under the following conditions which are critical to global $C^{1,1}$ estimates
(1.6) $\Delta(f^{1/(n-1)})\geq-A,\quad|D(f^{1/(n-1)})|\leq A,$
where $A$ is a constant. Guan [13] found that conditions (1.6) are also sufficient for the Dirichlet problem of degenerate Monge-Ampère equations with homogenous boundary. The non-homogeneous boundary case was resolved by Guan-Trudinger-Wang [14] under the stronger assumption $f^{\frac{1}{n-1}}\in C^{1,1}(\overline{\Omega})$, which implicitly enforces (1.6). For the Dirichlet problem of degenerate Hessian equations, Dong [6] established $C^{1,1}$ estimates to homogeneous boundary under conditions
$f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega}),\quad|D(f^{\frac{1}{k-1}})|\leq Cf^{\frac{1}{2(k-1)}}.$
Recently, $C^{1,1}$ estimates to convex solutions with non-homogenous boundary in the strictly convex bounded domain was solved by Jiao-Wang [23], requiring the condition $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$. For more references, the reader may see [2, 12, 46, 33] and related works cited there.
@@
Therefore, it is of natural interest to consider existence results for the oblique boundary value problem. The oblique boundary value problem for Hessian type equations arises naturally in the theory of fully nonlinear elliptic equations, through its applications in conformal geometry, optimal transportation, and geometric optics. Wang [44] established the existence of solutions for two dimensional Monge-Ampère equations and subsequently considered general dimensions. Urbas [41, 42, 43] studied Monge-Ampère equations, two dimensional Hessian equations and curvature equations of the oblique boundary value problem. The oblique boundary value problem of linear and quasilinear elliptic equations can be seen in the book [29]. Lieberman-Trudinger [30] investigated the nonlinear oblique boundary value problem for quasilinear and fully nonlinear uniformly elliptic partial differential equations under some structure conditions. For augmented Hessian equations with the oblique boundary value problem
@@
Following these developments, we investigate the oblique boundary value problem for degenerate Hessian quotient type equations under the assumption of subsolutions $\underline{u}$ with $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$, $0\leq l<k<n$ and $\gamma\geq 1$.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded $C^{4}$ domain, $\varphi(x,u)$ be smooth functions, $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$, $f_{u}\geq 0$, $\varphi_{u}\geq\gamma_{0}>0$. If there exists a subsolution $\underline{u}\in C^{2}(\overline{\Omega})$ satisfying
(1.7) $&\frac{\sigma_{k}}{\sigma_{l}}(\gamma\triangle\underline{u}I-D^{2}\underline{u})\geq f(x,\underline{u}),\quad\text{in}\,\;\Omega,\\ &\underline{u}_{\beta}=\varphi(x,\underline{u}),\quad\quad\quad\quad\quad\quad\quad\,\,\text{on}\,\;\partial\Omega,\end{aligned}\right.$
and a unique admissible solution $u\in C^{4}(\overline{\Omega})$ satisfying
(1.8) $&\frac{\sigma_{k}}{\sigma_{l}}(\gamma\triangle uI-D^{2}u)=f(x,u),\quad\text{in}\,\;\Omega,\\ &u_{\beta}=\varphi(x,u),\quad\quad\quad\quad\quad\quad\quad\,\,\text{on}\,\;\partial\Omega,\end{aligned}\right.$
where $\langle\beta,\nu\rangle\geq\beta_{0}>0$, $\gamma\geq 1$, $0\leq l<k<n$. Then
@@
where the positive constant $C$ depends on $n,\ k,\ l,\ \gamma_{0},\ \gamma,\ \beta_{0},\ \overline{\Omega}$, $G$, $|\underline{u}|_{C^{0}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$.
@@
The admissible subsolution hypothesis and conditions $f_{u}\geq 0$, $\varphi_{u}\geq\gamma_{0}>0$ are required only for the maximum modulus estimates. Furtherly, When $\gamma>1$ in problem (1.8), the existence theorem extends to the case $k=n$.
@@
When the matrix $\left\{\gamma\triangle uI-D^{2}u\right\}$ ($\gamma\geq 1$) is replaced by $\left\{\gamma\triangle uI+D^{2}u\right\}$ ($\gamma>0$) in problem (1.8), the existence theorem remains valid for the problem
(1.10) $&\frac{\sigma_{k}}{\sigma_{l}}(\gamma\triangle uI+D^{2}u)=f(x,u),\,\,\,\text{in}\,\;\Omega,\\ &u_{\beta}=\varphi(x,u),\quad\quad\quad\quad\quad\quad\quad\,\text{on}\,\;\partial\Omega,\end{aligned}\right.$
@@
The remainder of this paper is organized as follows. In Section 2, we introduce basic notations and establish preliminary results for Hessian quotient type equations. Section 3 is devoted to deriving $C^{0}$ estimates under the assumption of admissible subsolutions with $f_{u}\geq 0$ and $\varphi_{u}\geq\gamma_{0}>0$, followed by $C^{1}$ estimates under semilinear oblique boundary conditions. Finally, Section 4 presents the crucial second order estimates that complete the proof of Theorem 1.3.
@@ 2. Preliminaries
In this section, we recall some basic properties of Hessian quotient type equations, which can be found in [3, 5, 7, 31, 36]. For convenience, we define $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$, $\eta=(\eta_{1},\eta_{2},\cdots,\eta_{n})$ are eigenvalue vectors of matrix $D^{2}u$ and $U=\gamma\triangle uI-D^{2}u$, respectively, set
$F(u_{ij})=\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})},\quad F^{ij}=\frac{\partial F}{\partial u_{ij}},\quad F^{ij,rs}=\frac{\partial^{2}F}{\partial u_{ij}\partial u_{rs}},$
$\widetilde{F}(u_{ij})=\left[\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})}\right]^{\frac{1}{k-l}},\quad\widetilde{F}^{ij}=\frac{\partial\widetilde{F}}{\partial u_{ij}},\quad\widetilde{F}^{ij,rs}=\frac{\partial^{2}\widetilde{F}}{\partial u_{ij}\partial u_{rs}},$
@@
Differentiating equation (2.1) twice with respect to vector fields $\xi=(\xi^{1},\xi^{2},\cdots,\xi^{n})$, we have
(2.2) $F^{ij}u_{ij\xi}=\partial_{\xi}f=f_{x_{\xi}}+f_{z}u_{\xi},$
(2.3) $\widetilde{F}^{ij}u_{ij\xi}=\partial_{\xi}\widetilde{f}=\widetilde{f}_{x_{\xi}}+\widetilde{f}_{z}u_{\xi},$
@@
(2.4) $F^{ij,rs}u_{ij\xi}u_{rs\xi}+F^{ij}u_{ij\xi\xi}=\partial_{\xi\xi}f=f_{x_{\xi}x_{\xi}}+2f_{x_{\xi}z}u_{\xi}+f_{zz}u_{\xi}^{2}+f_{z}u_{\xi\xi},$
(2.5) $\widetilde{F}^{ij,rs}u_{ij\xi}u_{rs\xi}+\widetilde{F}^{ij}u_{ij\xi\xi}=\partial_{\xi\xi}\widetilde{f}=\widetilde{f}_{x_{\xi}x_{\xi}}+2\widetilde{f}_{x_{\xi}z}u_{\xi}+\widetilde{f}_{zz}u_{\xi}^{2}+\widetilde{f}_{z}u_{\xi\xi}.$
The following lemmas can be referenced to [5, 31, 36].
@@
If $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\Gamma_{k}$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ for $1\leq k\leq n$, then
@@
where $\sigma_{k-1}(\lambda|i)=\frac{\partial\sigma_{k}}{\partial\lambda_{i}}$ and $C(n,k)$ is a positive constant only depending on $n$ and $k$.
@@
If $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\widetilde{\Gamma}_{k}$, then $\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]^{\frac{1}{k-l}}(0\leq l<k\leq n)$ are concave with respect to $\lambda$. Hence, for any $\xi\in\mathbb{R}^{n}$
(2.6) $\sum_{i,j,r,s}\frac{\partial^{2}\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial u_{ij}\partial u_{rs}}\leq\left(1-\frac{1}{k-l}\right)\frac{\left[\sum\limits_{i,j}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial u_{ij}}\xi_{ij}\right]^{2}}{\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}},$
where $\widetilde{\Gamma}_{k}$ is an convex symmetric cone in $\mathbb{R}^{n}$, given by
@@
If $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\widetilde{\Gamma}_{k}$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, $\eta_{i}=\gamma\sum_{j=1}^{n}\lambda_{j}-\lambda_{i}$, $1\leq i\leq n$, for $0\leq l<k<n$ we obtain
@@
(3) $\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{i}}\geq\frac{2\gamma-1}{n\gamma-1}C(n,k,l)\sum_{j=1}^{n}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{j}},$
(4) $\sum_{j=1}^{n}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{j}}\geq C(n,k,l)f^{1-\frac{1}{k-l}}.$
@@
Compared to the classical Hessian equation with the internal matrix $D^{2}u$, the Hessian type equation involving $\gamma\triangle uI-D^{2}u$ exhibits uniform ellipticity which is indispensable for subsequent the a priori estimates.
Remark 2.6.
When the matrix $\left\{\gamma\triangle uI-D^{2}u\right\}$ ($\gamma\geq 1$) is replaced by $\left\{\gamma\triangle uI+D^{2}u\right\}$ ($\gamma>0$), the analogues of Lemma 2.3 and Lemma 2.4 still hold for $0\leq l<k\leq n$. Furthermore, the corresponding a priori estimates can be established through similar arguments, whose details we omit here.
3. $C^{0}$ and $C^{1}$ estimates
In this section, we first prove $C^{0}$ estimates for admissible solutions to problem (1.8) under the subsolution assumption with structural conditions $f_{u}\geq 0$, $\varphi_{u}\geq\gamma_{0}>0$, then proceed to derive $C^{1}$ estimates.